OpenStudy (hexagon001):

what would be the poles for the follwoing funciton? 1/(2z^2 +7iz − 3)^2 would it be pole at z=-i/2 of order 2 and pole at z=-3i of order 2??

OpenStudy (experimentx):

solve (2z^2 +7iz − 3) = 0

OpenStudy (experimentx):

both roots would be poles of it ... since it has ^2 ... should of order 2

OpenStudy (hexagon001):

but all that is also squared too... anyhow I get (2z+i)^2(z+3i)^2=0 so z=-i/2 of order 2 and z=-3i of order 2

OpenStudy (experimentx):

yes ...

OpenStudy (hexagon001):

thanks

OpenStudy (experimentx):

you should always follow from definition.